OpenStudy (anonymous):

write the equation of a line passing through (16,-1) and a slope of -1/4

hartnn (hartnn):

use the slope point form m = -1/4 x1 = 16 y1 =-1 in \(\large y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

you're welcome ^_^ so, what did you get the equation ? if you want me to verify..... oh and since you're new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{S}}\color{#e6e600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{ }}\color{#e6e600}{\text{:}}\color{green}{\text{)}}\color{blue}{\text{}}\end{array} \]

OpenStudy (anonymous):

I got y=-1/4x-5

hartnn (hartnn):

you sure about the -5 part ?

hartnn (hartnn):

m(-x1) = (-1/4)* (-16) = +4 so wouldn't it be y = -1/4 x + 3 ?

OpenStudy (anonymous):

but it's not -16...?

OpenStudy (anonymous):

oh wait never mind i get what i did wrong

hartnn (hartnn):

cool :)

OpenStudy (anonymous):

yea i got y=-1/4x=3 thanks for he help this place is cool

OpenStudy (anonymous):

*+

hartnn (hartnn):

this place is even cooler for good learners like you ;) if you have any doubts browsing this site, you can ask me :)

OpenStudy (anonymous):

thanks for the welcome, i'll be sure your the first to know of any complications :)

hartnn (hartnn):

welcome for the thanks :) lol